Mailings Review View Strong Subeitie 9-3 Stock Price vs. Intrinsic Value The intrinsic value of a stock is $39/share. The market value is curently $40/share. Should an investor buy or sell the stock? 9-4 The Dividend Discount Model How can the price of a stock be estimated? 9-5 Constant Growth Stocks-Part 1 A dividend grows from $1 in year 1, to $1.02 in year 2, to $1.0404 in year 3, and is expected to continue the same rate of growth forever. The required return on the stock is 8%. What is the price of the stock today? P-1/(0.08-0.02) $16.67 9-5 Constant Growth Stocks-Part 2 Todays price for a share of XYZ Corp. is $30/share. The dividend paid at the end of the first period is $3/share. At the end of that period, the stock price surged to $40/share. Calculate the total return on a share of XYZ Corp 9-6 Valuing Nopconstant Growth Stocks Part I Calculate the current price of a stock given the following information The required rate of return of the stock is 7%. From time period 0 to the end of the second period, the growth rate is 20%. From the beginning of time period 3 to the beginning of time penod 4, the growth rate is 5%. From the beginning of time period 4, the dividend grows 3% in perpetuity. Do $3.00 9-6 Valuing Nonconstant Growth Stocks-Part 2 Calculate the current price of a stock given the following information A dividend hasa growth rate of 0% for the first two p nodsBeginning period 3 and continuing in perpetuity, the dividend grows at 4%. The required rate ofretum is 7%. D-S200
Answer #1

As per rules I am answering the first 4 subparts of the question

9-3 The stock should be sold since the investor will get higher amount than its intrinsic value.

9-4: Price of stock is the present value of the dividends plus the horizon value of the stock.

9-5:

Growth rate = (d1-D0)/d0 = (1.02-1)/1 = 2%

In year 2, growth rate = (d2-d1)/d1 = (1.0404-1.02)/1.02= 2%

Hence price= D1/(k-g)

= 1/(8%-2%)

=16.67

9-5- Part 2

Total return= (P1-P0+dividend)/ P0

= (40-30+3)/30

=43.33%